OpenStudy (anonymous):

Members of a senior class held a car wash to raise funds for their senior prom. they charged $3 to wash a car and $5 to wash a pick up truck. If they earned a total of $275 by washing a total of 75 vehicles, how many cars did they wash?

OpenStudy (anonymous):

would the equation be something like 3c+5(75-c)=275 ?

OpenStudy (asapbleh):

3c+5t=275 c+t=75 t=75-c 3c+5(75-c)=275 3c+375-5c=275 -2c=-100 c=50 try to plug it in now

OpenStudy (anonymous):

okay thanks :) i got 50 too.. just cheking :)

OpenStudy (asapbleh):

kk